OpenStudy (bossimbacon):

The perimeter of a rectangle is less than 70 ft. The length of the rectangle is 5 ft longer than its width. Write an equation representing the length (L = length) of the rectangle in terms of Its width. L = _____________ b) The formula for perimeter of a rectangle is 2W + 2L = P W is the width, L is the length, and P is the perimeter. Write an inequality using the formula (plug in the value of L found in part a, the value of P, and replace the equal sign with the appropriate inequality

OpenStudy (bossimbacon):

@uri

OpenStudy (yttrium):

This is easy. Knowing that Perimeter is 70ft, we can say that P = 70 Also, given that the length of the rectangle is 5 ft more that its width, we can say that W = L - 5, where L and W represents length and width, respectively. Due to the situation given in the problem, length will still be represented by L. You know that P = 2L + 2W, you can now make an equation of L in terms of its width.

OpenStudy (bossimbacon):

@yttrium huh

OpenStudy (bossimbacon):

@Yttrium could you say that again

OpenStudy (yttrium):

Tell what you didn't get.

OpenStudy (bossimbacon):

the whole process could you show the answer and show me how you got there?

OpenStudy (bossimbacon):

@Yttrium

OpenStudy (yttrium):

That is not valid here. We don't give the fish. We teach you how to fish.

OpenStudy (bossimbacon):

@Yttrium thats why could you teach me?

OpenStudy (yttrium):

First, take note that P = 2L + 2W

OpenStudy (bossimbacon):

ok

OpenStudy (bossimbacon):

so if you said that w = L-5 then does that mean it becomes 70 = 2L + 2L-5

OpenStudy (bossimbacon):

?

OpenStudy (yttrium):

That is right. And you can now solve for the value of L

OpenStudy (bossimbacon):

do i add 5 to both sides of the equation to get rid of the 5 and then divide 2 from both sides of the equation to get rid of 2? @Yttrium

OpenStudy (yttrium):

You missed one step! That is 2L+2L = 4L

OpenStudy (bossimbacon):

ohhh so it would be 70 = 4L-5?

OpenStudy (bossimbacon):

and then from here I subtract 5 from both sides.

OpenStudy (yttrium):

Actually, add 5. And divide both sides by 4.

OpenStudy (bossimbacon):

oh yeah sorry and that would = 18.75 am I right? @ yttrium

OpenStudy (yttrium):

Correct!

OpenStudy (yttrium):

Let me take this as your bonus. P = 2L + 2W P - 2W = 2L (P - 2W)/2 = L, P = 70 L = (70-2W)/2 L = 35 - W ---> That's the equation of the length in terms of width

OpenStudy (bossimbacon):

but does it answer the queston>

OpenStudy (bossimbacon):

b) The formula for perimeter of a rectangle is 2W + 2L = P W is the width, L is the length, and P is the perimeter. Write an inequality using the formula (plug in the value of L found in part a, the value of P, and replace the equal sign with the appropriate inequality symbol). does it answer this question @Yttrium

OpenStudy (bossimbacon):

and did we find the length?

OpenStudy (bossimbacon):

was the length 18.75

OpenStudy (yttrium):

Sorry I need to go now buddy. Maybe next time

OpenStudy (bossimbacon):

@robtobey